OpenStudy (anonymous):

Using the Pythagorean Theorem, find the length of a leg of a right triangle if the other leg is 8 feet long and the hypotenuse is 10 feet long.

OpenStudy (anonymous):

\[a^2+b^2=c^2 \] \[a^2+8^2=10\] \[a^2+64=100\] \[a^2=36\ \[\sqrt{a}=\sqrt{36}\] A=6 c: